# Mastering Collaborative Multi-modal Data Selection: A Focus on Informativeness, Uniqueness, and Representativeness

### v1

@@ B.1 Cross-modal Domain Clustering
where $\sigma_{0}$ is the largest singular value of the feature matrix of each sample and $\sigma_{j}$ is the other singular value of the feature matrix. To amplify the contribution of task difficulty to data selection, we square the average maximum singular value ratio and normalize it based on the number of samples corresponding to each task, yielding the data selection rate as follows:
@@ C.1 Implemental Details
During data selection in DataTailor, we first normalize the uniqueness values across different clusters by dividing the intra-cluster uniqueness values by the average distance within each cluster, resulting in the unified uniqueness value. After unifying the principled values across samples, we enable collaboration among Informativeness, Uniqueness, and Representativeness values. Specifically, within the same task, we uniformly scale the values of informativeness, uniqueness, and representativeness to the range of [0, 1], ensuring consistency in their distributions. This approach facilitates balanced collaboration among these metrics in our adaptively collaborative data selection.
@@ C.2 Candidate Datasets Details
LLaVA-v1.5-mix-665k. This is currently the most extensive multimodal instruction dataset, encompassing instruction data across a wide range of tasks. It contains a variety of datasets: VQA [5], OCR [38], region-level VQA [22], visual conversation [35] and language conversation [1] data. For all datasets, QA pairs from the same training image are merged into a single conversation, and excessively long data is filtered out to improve training efficiency. As a result, this process yields 665k instruction pairs across 10 tasks.
@@
Methods Data Ratio LLaVA-Wild MM-Vet
LLaVA-v1.5-7B (LoRA) [34] 100% 84.3 30.9
   Random 7.5% 82.6 29.5
   Length 7.5% 84.5 29.7
   E2LN [42] 7.5% 40.1 21.1
   GradN [42] 7.5% 68.9 24.8
   IFD [30] 7.5% 81.9 27.6
   InsTag [37] 7.5% 84.4 29.6
   LESS [54] 7.5% 83.1 28.3
   SELF-FILTER [53] 7.5% 80.5 26.6
   TIVE [36] 7.5% 84.3 30.2
   DataTailor (Ours) 7.5% 85.0 30.4
   DataTailor w/ Increased Ratio (Ours) 15% 85.9 31.8
Table 7: Open-ended evaluation of data selection methods to show their robustness for real-world application. Bold and underline fonts indicate the best and second-best performance on the task.
@@ Appendix D Additional Experiment Analyses
D.1 Open-ended Evaluation
To verify that high-quality data selected from DataTailor effectively supports the open-ended capabilities of MLLMs for wide applications, we further compare MLLMs fine-tuned on DataTailor-selected data with baselines on open-ended benchmarks (i.e., LLaVA-Wild [35] and MMVet [64]) in Table 7. Following prior works [35, 34], We prompt GPT-4 to compare the answers generated by MLLM with those produced by text-only GPT-4, providing a rating and an accompanying explanation. We observe that DataTailor achieves promising results on these open-ended questions, delivering competitive performance with only 7.5% of the data (85.0 on LLaVA-Wild and 30.4 on MM-Vet). As the data selection ratio increases, our DataTailor significantly outperforms fine-tuning on the full dataset, achieving 85.9 vs. 84.3 on LLaVA-Wild and 31.8 vs. 30.9 on MM-Vet. It demonstrates that the high-quality data selected by DataTailor, based on three principles, not only retains discriminative capabilities but also enhances generative abilities, promoting the open-ended responses required in real-world MLLM applications.
D.2 Human Evaluation

### v2

@@ B.2 Adaptive Data Proportion for Data Selection
To amplify the contribution of task difficulty to data selection, we square the average largest singular value ratio and normalize it based on the number of samples corresponding to each task, yielding the data selection ratio as follows:
@@
B.3 Cross-Modal Domain Clustering
B.3.1 Algorithmic Formulation
The clustering pipeline commences with inter-sample affinity quantification via $\ell_{2}$-norm distance measurements within each initial category. Starting with all nodes as individual clusters, we iteratively merge the pair (A,B) exhibiting the minimal increase in Sum of Squared Errors (SSE) to construct a dendrogram:
$A,B=\mathop{\arg\min}\limits_{A,B\in\mathcal{P}}\Delta\mathrm{SSE}(A,B)$ (12)
where $\mathcal{P}$ denotes the partition of nodes into distinct clusters. The SSE increment from merging clusters $A$ and $B$ is defined as:
$\Delta\mathrm{SSE}(A,B)=\frac{n_{A}n_{B}}{n_{A}+n_{B}}\|\bm{\mu}_{A}-\bm{\mu}_{B}\|_{2}$ (13)
with $n_{A},n_{B}$ representing cluster sizes and $\bm{\mu}_{A},\bm{\mu}_{B}$ their centroids. This criterion minimizes intra-cluster variance growth. To theoretically characterize the merging behavior, we formalize a key monotonicity property inherent to Ward’s algorithm.
Theorem 1 (SSE Monotonicity in Ward’s Method).
For a hierarchical merging process $\{\mathcal{P}_{k}\}_{k=0}^{n-1}$ under Ward’s algorithm (where $\mathcal{P}_{0}$ contains singleton clusters), the total SSE increment satisfies the non-decreasing property:
$\Delta\mathrm{SSE}(\mathcal{P}_{k})\geq\Delta\mathrm{SSE}(\mathcal{P}_{k-1}),\quad\forall k\geq 1,$
where $\Delta\mathrm{SSE}(\mathcal{P}_{k})$ denotes the incremental SSE from merging $\mathcal{P}_{k-1}$ to $\mathcal{P}_{k}$.
This monotonic progression ensures that earlier merges correspond to more natural cluster unions, while later merges sacrifice increasing amounts of variance. Leveraging this property, we define a threshold-based partitioning rule to extract clusters from the dendrogram hierarchy. A dendrogram $T=(V,E)$ consists of leaf nodes $V_{\text{leaf}}=\{s_{1},...,s_{n}\}$ and internal nodes, which serve as merge points annotated with $\Delta\mathrm{SSE}$ values. Given a threshold $T=\lambda\cdot\Delta\mathrm{SSE}(\mathcal{P}_{n-1})$, the optimal partition is determined by:
$\mathcal{P}^{*}=\max\left\{k\mid\Delta\mathrm{SSE}(\mathcal{P}_{k})\leq T\right\}$
B.3.2 GPU-Accelerated Distance Computation
For efficient pairwise distance measurement between $n$ samples in $\mathbb{R}^{d}$, we implement a parallelized $\ell_{2}$-norm computation framework using CUDA-optimized matrix operations. The distance matrix $D\in\mathbb{R}^{n\times n}$ can be derived through the algebraic identity, which:
$D_{x,y}=\sqrt{S_{x,x}+S_{y,y}-2S_{x,y}}$ (14)
where $X\in\mathbb{R}^{n\times d}$ represents the feature matrix containing $n$ samples, and $S=XX^{T}\in\mathbb{R}^{n\times n}$ represents the similarity matrix. Our kernel-based implementation achieves $\mathcal{O}(n^{2}d/m)$ theoretical speedup through massive parallelization across GPU cores, effectively transforming an originally $\mathcal{O}(n^{2}d)$ complexity operation into a highly parallelizable matrix multiplication task. In addition, we use the parallel pipeline strategy to extract feature vectors of the feature matrix while calculating the $\ell_{2}$-norm distance between sample features for efficiency.
B.3.3 Optimized Hierarchical Clustering Merge
To address the quadratic complexity inherent in conventional hierarchical clustering, we implement a memory-efficient variant of the nearest-neighbor chain algorithm. This optimization framework features:
  • Randomized stack initialization with cluster prototypes
  • Iterative nearest-pair identification via stack
  • In-stack merging with $\mathcal{O}(n)$ per-operation cost, where each node undergoes $\mathcal{O}(1)$ amortized operations.
The proposed acceleration strategy reduces computational overhead to $\mathcal{O}(n^{2})$ while preserving the theoretical guarantees of Ward’s method. By optimizing the clustering process, the time consumption of the clustering was accelerated by $90\times$, resulting in significantly improved performance.
Threshold $\lambda$ w/o clustering 0.05 0.1 0.25 0.5
MLLM Rel. 98.1% 99.8% 101.3% 100.0% 98.5%
B.3.4 Threshold Parameter Analysis
Since the quality of clustering is critical for the domain-based adaptive data proportion in DataTailor, we further explore the effect of dynamic threshold to cross-modal domain clustering on the multi-modal data selection in Table 7. The clustering threshold determines the cluster size based on the dataset’s sample distribution. Therefore, setting it close to the overall data selection proportion ensures an appropriate size of clusters that effectively captures sample relationships. Our experiments reveal that low or high thresholds compromise the constraints on the uniqueness or representativeness of high-quality samples, leading to lower performance of DataTailor. Thus, we set the appropriate threshold $\lambda$ as $0.1$ for cross-modal domain clustering, which is close to the total data selection proportion.
B.4 Balance between Three Principles
Since multi-modal samples exhibit varying structures, we propose an adaptive weight to combine the three principal values. We restate the underlying principles behind the three properties to show their effectiveness: (a) Informativeness. It determines external relationships due to its core training contributions. Because MLLMs rely on token-level inputs, the SVD of token feature space reveals the samples’ contribution to the MLLM. That’s why we use singular value entropy to reflect the value of samples for generalization. (b) Uniqueness. For repeated samples, their uniqueness is adjusted during selection based on the distribution of chosen points by normalization, ensuring duplicates are treated differently. (c) Representativeness. It aims to isolate undesired noisy samples that may have a high uniqueness score, while general noisy data can be identified by its lower informativeness (i.e., average 0.297 in clusters).
In general, it is important to balance the above three principles to select data collaboratively. On the one hand, we explore various ratios of these two external values (i.e., uniqueness and representativeness) in Figure 8. The smooth performance transition (less than 2%) near the 1:1 ratio suggests that the trade-off between them remains stable. Therefore, we select the 1:1 ratio as the optimal value for collaboration. On the other hand, we use adaptive weights between the information values and the two values for the varying instruction rounds of the samples in the dataset.
$V^{Uni}:V^{Rep}$ 0:1 0.5:1 1:1 1:0.5 1:0
MLLM Rel. 99.6% 100.5% 101.3% 100.3% 99.5%
Table 8: Balance between uniqueness and representation in DataTailor for data selection of MLLMs.
@@ C.1 Implementation Details
During data selection in DataTailor, we balance the uniqueness and representativeness of different clusters. First, we normalize uniqueness and representativeness values across clusters by removing the influence of spatial distribution and cluster size. Next, we standardize the impact of the average sample value across clusters. Specifically, within the same task, we uniformly scale informativeness, uniqueness, and representativeness metrics to a $[0,1]$ range to ensure consistent distribution alignment. By harmonizing these normalized values across samples, we enable collaboration among Informativeness, Uniqueness, and Representativeness values. This methodology fosters balanced metric collaboration in our adaptive data selection framework, ensuring proportional consideration of all three criteria.
@@
Methods Selected Data Evaluation MLLM Rel.
Informativeness Uniqueness Representativeness
IFD (7.5%) 32.3 0.341 30.3 87.3%
INSTAG (7.5%) 30.9 0.347 34.4 96.4%
LESS (7.5%) 34.0 0.314 33.3 94.3%
   $V_{i}^{Inf}$ (7.5%) 34.5 0.348 34.8 98.0%
   $V_{i}^{Uni}$ (7.5%) 33.4 0.364 34.5 97.3%
   $V_{i}^{Rep}$ (7.5%) 33.9 0.343 35.0 97.5%
DataTailor (7.5%) 34.8 0.358 34.9 100.1%
Table 9: Quantitative Valuation of Three Principles. All setups are on the model of LLaVA-v1.5-7B and the dataset of LLaVA-mix-665K for both data selection and MLLM training.
@@ C.2 Candidate Datasets Details
LLaVA-v1.5-mix-665k. This is currently the most extensive multimodal instruction dataset, encompassing instruction data across a wide range of tasks. It contains a variety of datasets: VQA [5], OCR [45], region-level VQA [28], visual conversation [42] and language conversation [1] data. For all datasets, QA pairs from the same training image are merged into a single conversation, and excessively long data is filtered out to improve training efficiency. As a result, this process yields 665k instruction pairs across 13 tasks.
@@ Appendix D Additional Experiment Analyses
D.1 Quantitative Valuation of Three Principles
To quantitatively investigate how DataTailor addresses each of the three principles for data selection, we designed three experimental settings to assess them. Here we show the detailed metric values for DataTailor and other baselines on three experimental setups in Table 9.
Methods Dataset MMMU MathVerse MathVision MLLM Avg.
QVQ-72B-preview - 66.0 41.5 38.2 48.6
InternVL2.5-78B - 70.0 39.2 32.2 47.1
Qwen2-VL-7B Virgo(100%) 46.7 36.7 24.0 35.8
   w/ random Virgo(15%) 46.7 32.5 23.4 34.2
   w/ DataTailor Virgo(15%) 50.1 34.8 23.8 36.2
Qwen2-VL-72B Virgo(100%) 65.0 48.1 38.6 50.6
   w/ random Virgo(15%) 59.4 43.7 36.8 46.6
   w/ DataTailor Virgo(15%) 63.0 46.2 40.3 49.8
Table 10: The scalable results on math and reasoning benchmarks of DataTailor with more scalable MLLMs.
D.2 More scalable MLLMs Results
Further, we apply DataTailor to larger and newer backbones (i.e., Qwen-2-VL-7B & 72B) for more robust evaluation in the Table 10. Since Qwen’s data is closed-source, we fine-tune on the open Virgo dataset [15]. Similarly, we observe that DataTailor exhibits competitive performance with only 15% data (36.2 v.s. 35.8 of full data), outperforming other open-source MLLMs baselines.
D.3 Instantiation Comparisons of Three Principles
To fully explore the characteristics of valuable samples that are meaningful for downstream tasks, we further present a few instantiation comparisons of sample characteristics from the three principal perspectives in Figure 8. The left side shows the data preferred by DataTailor, and the right side shows the remaining data.
From the perspective of informativeness, we can observe that the samples selected by DataTailor contain richer information and various description, whereas other samples suffer from excessive redundancy in responses and numerous blank pixels in images (e.g., the right image shows only snow slope and the text repeating snowboard). This indicates that selecting samples according to informativeness can select samples as complex as possible to improve inference accuracy in downstream tasks while also facilitating the generation of richer and higher-quality content.
From the perspective of uniqueness, we can observe that the samples selected by DataTailor in each cluster contain unique insights (e.g., critically mention artistic installation) and in-depth novel analysis (e.g., analyze diversity of the ornamental tree). However, other examples in the cluster show similar information and primarily describe basic, pre-learned commonsense knowledge, which limits MLLMs’ ability to enhance generalization on downstream tasks continuously. This suggests that selecting samples based on uniqueness enables the inclusion of distinctive samples, which provide deeper and more novel insights to continuously enhance the reasoning capability during the fine-tuning phase of MLLM.
From the perspective of representativeness, we observe that the samples selected by DataTailor exhibit accurate descriptions and aligned answers. This is because the method ensures that the selected samples represent the overall distribution, avoiding noise and mislabeled data. Samples with low representativeness tend to exhibit mislabeled answers and outlier features, which can lead to incorrect optimization directions, ultimately hindering the performance improvement of downstream tasks. This suggests the need for assessing the representativeness of samples to filter out noisy or mislabeled data.
Methods Informativeness Uniqueness Representativeness MLLM Rel.
$V^{Inf}$ (15%) 34.3(33.1) - - 99.2%(97.7%)
$V^{Uni}$ (15%) - 0.363(0.347) - 98.5%(97.4%)
$V^{Rep}$ (15%) - - 34.6(32.5) 98.6%(97.6%)
Table 11: Deep Analysis for Calculation of Three Values
Figure 8: Instantiation Comparisons of DataTailor in addressing three core principles (i.e. informativeness, uniqueness, and representativeness) for selecting valuable multi-modal instruction samples.
Based on the above, we propose three essential principles (i.e., informativeness, uniqueness, and representativeness) for the practical selection of valuable samples that truly contribute to the downstream inference performance of MLLMs. To further illustrate the superiority of these three quantitative metrics for three principles, we explore the insights underlying three values and conduct more ablation of potential alternatives (results are of red color in Table 11) For informativeness, we adopt the SVE based on the theory that more informative samples have feature matrices closer to full rank, leading to more uniform singular values and thus higher SVE. We explore pair-wisely calculating mutual information between token-level features as an alternative. As SVE captures the overall distribution of feature directions, it achieves better informativeness and MLLM performance for data selection (99.2% v.s. 97.7%). For Uniqueness, we adopt the Euclidean distance based on the theory that larger distances in the feature space separate more unique samples and samples containing pre-learned common knowledge are clustered near the center. We explore directly selecting one sample from each cluster as an alternative. As Euclidean distance measures samples’ deviation from dense regions, it obtains better MLLM performance with uniqueness (98.5% v.s. 97.4%). For Representativeness, we adopt cosine similarity based on the theory that it reflects the true directional alignment of the overall distribution. We explore using the overall Euclidean distance between samples as an alternative. As cosine similarity emphasizes overall directional alignment instead of being influenced by the magnitude of outlier samples like distance, it performs better in MLLM with true representativeness (98.6% v.s. 97.6%).
D.4 Limitations
We observe some failure cases that DataTailor cannot discriminate long-term instructions when the reasoning process is omitted, e.g., math reasoning and code program.
D.5 Human Evaluation